# Optimal Stopping for Sequential Bayesian Experimental Design

### v1

@@ 4.2 Policy gradient for optimal stopping
Theorem 1 shows that the optimal stopping policy is implicitly determined by the design policy $\pi$. This suggests that the joint optimization problem can be reformulated as a single optimization over the design policy alone. We therefore parameterize the design policy, derive the gradient of the expected utility with respect to its parameters, and employ gradient-based optimization. This leads to a policy gradient (PG) method (Silver et al., 2014; Wang et al., 2020).
@@ 4.2.1 Policy gradient derivation
$\mathcal{T}_{k,w}$ $=\left\{s_{k}\,\Big|\,r^{S}_{T}(s_{k})\geq\mathbb{E}_{y_{k}\mid\pi_{w},s_{k}}\Big[r^{C}_{k}(s_{k},\mu_{k,w_{k}}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}\big(\mathcal{F}_{k}(s_{k},\mu_{k,w_{k}}(s_{k}),y_{k})\big)\Big]\right\}$
@@
$U(w)$ $=\mathbb{E}_{y_{0:N-1}\mid\pi_{w},\psi_{w},s_{0}}\Bigg[\sum_{k=0}^{N-1}\left(r_{k}(s_{k},\mu_{k,w_{k}}(s_{k}),y_{k})+r_{T}(s_{k})\right)+r_{T}(s_{N})\Bigg].$ (17)
Theorem 3 (Policy Gradient).
@@
$\nabla_{w}U(w)=\sum_{k=0}^{N-1}\mathbb{E}_{s_{k}\mid\pi_{w},\psi_{w},s_{0}}\left[\bm{1}_{k<\tau_{w}}\nabla_{w}\mu_{k,w_{k}}(s_{k})\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w_{k}}(s_{k})}\right],$ (18)
@@ 4.2.2 Numerical estimation
The gradient in Theorem 3 involves nested expectations over stochastic trajectories and stopping decisions, which are intractable in closed form. We therefore use a Monte Carlo (MC) estimator:
$\nabla_{w}U(w)\approx$ $\frac{1}{M}\sum_{m=1}^{M}\sum_{k=0}^{N-1}\bm{1}_{k<\tau^{(m)}}\nabla_{w}\mu_{k,w_{k}}\left(s^{(m)}_{k}\right)\nabla_{\xi^{(m)}_{k}}Q^{\pi_{w},\psi_{w}}_{k}\left(s^{(m)}_{k},\xi^{(m)}_{k}\right)\Big|_{\xi^{(m)}_{k}=\mu_{k,w_{k}}\left(s^{(m)}_{k}\right)}$, (19)
@@
Two challenges arise: (1) policy gradients $\nabla_{w}\mu_{k,w_{k}}$ must be computed efficiently, and (2) Q-function gradients $\nabla_{\xi}Q^{\pi_{w},\psi_{w}}_{k}$ cannot be estimated reliably from nested MC sampling. We address both using neural network approximations in an actor-critic framework (Sutton and Barto, 2018).
@@ 4.2.3 Actor-Critic Implementation
Policy network. The design policy is parameterized by a single deep neural network $\mu_{w}(k,s_{k})$. Gradients $\nabla_{w}\mu_{k,w_{k}}\left(s^{(m)}_{k}\right)=\nabla_{w}\mu_{w}\left(k,s^{(m)}_{k}\right)$ are computed via standard backpropagation.
Q-network. To approximate the continuation value, we train a neural network $Q_{\eta}^{\pi_{w},\psi_{w}}(k,s_{k},\xi_{k})$ with parameters $\eta$. This avoids costly inner MC sampling for directly estimating $Q$ and provides differentiable Q-function estimates. The Q-network is trained by minimizing:
$\mathcal{L}(\eta)=$ $\frac{1}{M}\sum_{m=1}^{M}\sum_{k=0}^{N-1}\Bigg[Q^{\pi_{w},\psi_{w}}_{\eta}\left(k,s^{(m)}_{k},\xi^{(m)}_{k}\right)-\left(r_{k}\left(s^{(m)}_{k},\xi^{(m)}_{k},y^{(m)}_{k}\right)+Q^{\pi_{w},\psi_{w}}_{k+1}\left(s^{(m)}_{k+1},\xi^{(m)}_{k+1}\right)\right)\Bigg]^{2}\bm{1}_{k<\tau^{(m)}},$ (20)
@@
Training. We follow a standard actor-critic loop in Algorithm 1. During training, a parameter sample $\theta^{(m)}\sim p(\theta)$ generates an episode, and exploration is encouraged by injecting Gaussian noise to the design:
$\xi_{k}=\mu_{k}(s_{k})+\epsilon_{\text{explore}},\quad\epsilon_{\text{explore}}\sim\mathcal{N}(0,\mathbb{I}_{N_{\xi}}\sigma^{2}_{\text{explore}}).$ (21)
Stopping is determined when $r_{T}^{S}(s_{k})$ exceeds the continuation value $Q_{k}^{\pi,\psi}(s_{k},\mu_{k}(s_{k}))$ in (14). Policy parameters $w$ are updated via gradient ascent using the MC gradient estimator in (19), while Q-network parameters $\eta$ are updated via stochastic gradient descent on the loss in (20). See Appendix A.4 for more implementation details.
Algorithm 1 presents the overall pseudocode of our method.
@@
1: Set initial state $s_{0}$, policy updates $L$, MC sample size $M$, policy and Q-network architectures, learning rate $\alpha$ for policy update, exploration scale $\sigma_{\rm{explore}}$;
@@
7:   Sample design $\xi_{k}^{(m)}=\mu_{w}(k,s_{k}^{(m)})+\epsilon_{\text{explore}}$ and $y_{k}^{(m)}\sim p(\cdot\big|\theta^{(m)},\xi_{k}^{(m)},I_{k}^{(m)})$;
8:   Update state $s_{k+1}^{(m)}$, and compare $r_{T}^{S}(s^{(m)}_{k+1})$ with $Q_{\eta}^{\pi,\psi}(k+1,s^{(m)}_{k+1},\mu_{w}(k+1,s^{(m)}_{k+1}))$ for stopping decision;
@@ 4.3 Training instabilities and curriculum learning
Although the PG method provides a principled framework, implementing it in practice reveals training instabilities arising from the problem’s inherent circular dependency. The optimal stopping set $\mathcal{T}_{k}$ in (14) depends on the Q-function $Q_{k}^{\pi,\psi}(s_{k},\mu_{k}(s_{k}))$, which itself depends on both the design policy $\pi$ and stopping policy $\psi$. Thus, the stopping policy requires knowledge of the Q-function computed under that very same policy, leading to a fixed-point relationship $\psi^{*}=f(\pi,\psi^{*})$.
@@
In practice, this circularity manifests during joint training of the policy network (parameters $w$) and the Q-network (parameters $\eta$). Early in training, poorly initialized design policies generate weak experiments with little information gain. The Q-network then learns pessimistic continuation values, which in turn induce premature stopping that prevents policy improvement and entering suboptimal designs. This issue is particularly pronounced in sequential design problems with strong sequential dependencies, where early stopping disrupts the learning of effective long-term strategies.
To overcome this challenge, we introduce a curriculum learning scheme based on a stopping probability schedule $p_{\text{stop}}(\ell)$, where $\ell$ indexes training iterations. When the stopping condition $s_{k}\in\mathcal{T}_{k}$ is satisfied, the algorithm follows the optimal stopping rule with probability $p_{\text{stop}}(\ell)$, and overrides it (forcing continuation) with probability $(1-p_{\text{stop}}(\ell))$.
@@
This strategy deliberately relaxes stopping early in training, producing longer trajectories that allow both the policy and Q-networks to improve before stopping decisions dominate. As training progresses, $p_{\text{stop}}(\ell)$ is gradually increased, so that stopping behavior converges to the optimal fixed point. In this way, curriculum learning breaks the self-reinforcing cycle of premature stopping and stabilizes training, particularly in settings with strong sequential dependencies.
@@ 5 Numerical Experiments
We demonstrate our approach on two numerical examples: a linear-Gaussian benchmark for validation and a sensor movement problem for contaminant source detection that highlights the benefit of curriculum learning. The curriculum learning schedule is provided in Appendix A.6.1.
@@ 5.1 Linear-Gaussian Benchmark
with prior $\theta\sim\mathcal{N}(0,3^{2})$. Experiments are constrained to designs $\xi_{k}\in[0.1,3]$. Conjugacy allows closed-form solutions for optimal policies (see Appendix A.5).
@@
For zero cost ($c_{k}=0$), the analytical solution shows that stopping is always optimal at the horizon $N$, since the KL divergence reward is non-decreasing. Figure 3 (left column) shows that our method converges to the maximum stage $N=3$, achieving the analytical optimum, while naïve threshold policies stop prematurely and underperform.
For negative costs, Figure 3 (middle and right columns) illustrate two cases:
  • An $N=3$ problem with $c_{k}=-0.5$, where both vanilla and curriculum methods converge to the optimal early stopping at stage 1, consistent with analytical results.
  • An $N=4$ problem with cost $c_{k}=-0.25$, where both methods achieve near-optimal performance but with distinct convergence patterns. The vanilla approach often underestimates the optimal stopping stage, becoming trapped below the analytical solution. In contrast, the curriculum method converges more reliably, exploring longer trajectories before settling at the correct stage.
These results illustrate that curriculum learning promotes more stable convergence by mitigating premature stopping during training. The small discrepancy from theory in the $N=4$ case reflects the flat reward landscape, where stopping at stages 1–3 yields nearly indistinguishable rewards. Even so, the PG method successfully navigates these subtle trade-offs, validating both the algorithm and the role of curriculum learning.
5.2 Contaminant Source Detection in Convection-Diffusion Field
We next consider a sensor-movement problem for contaminant source detection, which exhibits strong sequential dependencies where curriculum learning is essential. Mobile sensors must be strategically repositioned to locate an unknown pollution source, with each placement decision depending critically on previous locations and measurements. Contaminant transport is modeled by a convection-diffusion partial differential equation (PDE) in a two-dimensional square domain with a Gaussian source term (see Appendix A.6.2 for details). We present constant-cost experiments below; design-dependent cost cases are given in Appendix A.6.3.

### v2

@@ 3.2 Policy gradient for optimal stopping
Theorem 3.1 shows that the optimal stopping policy is implicitly determined by the design policy $\pi$. This suggests that the joint optimization problem can be reformulated as a single optimization over the design policy alone. We therefore parameterize the design policy, derive the gradient of the expected utility with respect to its parameters, and employ gradient-based optimization. This leads to a policy gradient (PG) method, which directly optimizes the design policy by differentiating the expected utility with respect to its parameters (Silver et al., 2014; Wang et al., 2020).
@@ 3.2.1 Policy gradient derivation
$\mathcal{T}_{k,w}$ $=\Big\{s_{k}\ \Big|\ r^{S}_{T}(s_{k})\geq\mathbb{E}_{y_{k}\mid\pi_{w},s_{k}}\Big[r^{C}_{k}(s_{k},\mu_{k,w_{k}}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}\big(\mathcal{F}_{k}(s_{k},\mu_{k,w_{k}}(s_{k}),y_{k})\big)\Big]\Big\}$
@@
$U(w)=\mathbb{E}_{y_{0:N-1}\mid\pi_{w},\psi_{w},s_{0}}\Bigg[\sum_{k=0}^{N-1}(r_{k}(s_{k},\mu_{k,w_{k}}(s_{k}),y_{k})+r_{T}(s_{k}))+r_{T}(s_{N})\Bigg].$ (17)
Theorem 3.3 (Policy Gradient).
@@
$\nabla_{w}U(w)=\sum_{k=0}^{N-1}\mathbb{E}_{s_{k}\mid\pi_{w},\psi_{w},s_{0}}\Big[\mathbf{1}_{k<\tau_{w}}\nabla_{w}\mu_{k,w_{k}}(s_{k})\cdot\nabla_{\xi_{k}}Q^{\pi_{w},\psi_{w}}_{k}(s_{k},\xi_{k})\big|_{\xi_{k}=\mu_{k,w_{k}}(s_{k})}\Big]$ (18)
@@ 3.2.2 Numerical estimation
The gradient in Theorem 3.3 involves nested expectations over stochastic trajectories and stopping decisions, which are intractable in closed form. We therefore use a Monte Carlo estimator:
$\nabla_{w}U(w)\approx\frac{1}{M}\sum_{m=1}^{M}\sum_{k=0}^{N-1}\mathbf{1}_{k<\tau^{(m)}}\nabla_{w}\mu_{k,w}\left(s^{(m)}_{k}\right)\cdot\nabla_{\xi^{(m)}_{k}}Q^{\pi_{w},\psi_{w}}_{k}\left(s^{(m)}_{k},\xi^{(m)}_{k}\right)\Big|_{\xi^{(m)}_{k}=\mu_{k,w}\left(s^{(m)}_{k}\right)}$ (19)
@@
Two computational challenges arise: policy gradients $\nabla_{w}\mu_{k,w_{k}}$ must be evaluated efficiently, and Q-function gradients $\nabla_{\xi}Q^{\pi_{w},\psi_{w}}_{k}$ cannot be estimated reliably by nested Monte Carlo because such estimates would have high variance and prohibitive cost. We address both challenges with an actor-critic approximation (Sutton and Barto, 2018). The actor represents the design policy and is updated using the policy gradient estimator in (19), while the critic approximates the continuation value and supplies differentiable, lower-variance gradient information for improving the actor.
@@ 3.2.3 Actor-Critic Implementation
Policy network. The design policy is parameterized by a single deep neural network $\mu_{w}(k,s_{k})$ that takes as input both the current stage index $k$ and state $s_{k}$. Since the history grows in length across stages, we use a fixed-size, zero-padded representation as in Shen and Huan (2023). The stage index is encoded as a one-hot vector,
$k\qquad\longrightarrow\qquad e_{k+1}=[0,\dots,0,\underbrace{1}_{(k+1)\text{th}},0,\dots,0]^{T},$ (20)
and the state $s_{k}$ is represented in a nonparametric manner with the experimental history $I_{k}$,
which is laid out sequentially and zero-padded (for experiments that have not happened yet) to a fixed dimension $(N-1)(N_{\xi}+N_{y})$. The complete input vector at stage $k$ is
$\bigl[\,\underbrace{e_{k+1}}_{N},\;\overbrace{\xi_{0}}^{N_{\xi}},\dots,\xi_{k-1},\underbrace{0,\dots,0}_{N_{\xi}(N-1-k)},\;\overbrace{y_{0}}^{N_{y}},\dots,y_{k-1},\underbrace{0,\dots,0}_{N_{y}(N-1-k)}\,\bigr]^{T},$ (22)
giving a total input dimension of $N+(N-1)(N_{\xi}+N_{y})$. The output is an $N_{\xi}$-dimensional vector representing the proposed design $\xi_{k}$. Gradients $\nabla_{w}\mu_{k,w_{k}}\left(s^{(m)}_{k}\right)=\nabla_{w}\mu_{w}\left(k,s^{(m)}_{k}\right)$ are computed via standard backpropagation.
Q-network. To approximate the continuation value defined in (13), we train a neural network $Q_{\eta}^{\pi_{w},\psi_{w}}(k,s_{k},\xi_{k})$ with parameters $\eta$. This avoids costly inner Monte Carlo sampling for directly estimating $Q$ and provides differentiable Q-function estimates. The Q-network shares the same input structure as the policy network but additionally takes the candidate design $\xi_{k}$, giving total input dimension $N+(N-1)(N_{\xi}+N_{y})+N_{\xi}$; the output is a scalar representing the expected continuation value. The Q-network is trained by minimizing:
$\mathcal{L}(\eta)=\frac{1}{M}\sum_{m=1}^{M}\sum_{k=0}^{N-1}\Bigg[Q^{\pi_{w},\psi_{w}}_{\eta}\bigl(k,s^{(m)}_{k},\xi^{(m)}_{k}\bigr)-r_{k}\bigl(s^{(m)}_{k},\xi^{(m)}_{k},y^{(m)}_{k}\bigr)-Q^{\pi_{w},\psi_{w}}_{k+1}\bigl(s^{(m)}_{k+1},\xi^{(m)}_{k+1}\bigr)\Bigg]^{2}\mathbf{1}_{k<\tau^{(m)}},$ (23)
@@
Training. We follow a standard actor-critic loop in Algorithm 1. During training, each episode is generated by first sampling a parameter $\theta^{(m)}\sim p(\theta)$. The design policy $\mu_{k}(s_{k})$ itself is deterministic; however, to encourage exploration during trajectory generation, we inject a small perturbation into the design when generating the Monte Carlo samples used for gradient estimation:
$\xi_{k}=\mu_{k}(s_{k})+\epsilon_{\text{explore}},\hskip 3.99994pt\epsilon_{\text{explore}}\sim\mathcal{N}(0,\mathbb{I}_{N_{\xi}}\sigma^{2}_{\text{explore}}).$ (24)
The value of $\sigma_{\text{explore}}$ reflects the degree of exploration and should be selected based on the problem context. A common practice is to set a large $\sigma_{\text{explore}}$ early in training and reduce it gradually. Stopping is triggered deterministically when $r_{T}^{S}(s_{k})$ exceeds the continuation value $Q_{k}^{\pi,\psi}(s_{k},\mu_{k}(s_{k}))$ in (13). Policy parameters $w$ are updated via gradient ascent using the Monte Carlo gradient estimator in (19), while Q-network parameters $\eta$ are updated via stochastic gradient descent on the loss in (23).
Information gain computation. The terminal reward calculation requires computing KL divergences between posterior distributions. For problems without analytical posteriors, we approximate these quantities by discretizing the parameter space $\theta$ on a regular grid and estimating posterior densities pointwise. Figure 3 demonstrates the accuracy of this grid-based approximation in the linear-Gaussian case, where we compare the approximated posterior against the analytical solution. The close agreement validates that grid-based computation provides sufficient accuracy for reward evaluation in our experiments.
Figure 3: Comparison of grid-based posterior approximation against the analytical Gaussian posterior in the linear-Gaussian benchmark.
In implementation, the choice between terminal and incremental formulations involves a computational tradeoff. In the terminal-reward formulation, all intermediate rewards are zero, providing sparse supervision to the critic. This can slow convergence early in training, as the Bellman loss (23) receives gradient signal only from episodes that reach the stopping decision. Conversely, the incremental formulation provides denser reward signals at each stage, which can accelerate learning. However, this comes at the cost of computing information gains in $r_{k}(\cdot)$ at every stage, which is computationally expensive for high-dimensional posteriors and introduces approximation errors in intermediate posterior updates. In practice, the choice depends on the available computational budget and desired learning dynamics. In our experiments, we adopt the terminal formulation for its simplicity and transparent interpretation of the cumulative reward.
Algorithm 1 summarizes the resulting training procedure.
@@
1: Set initial state $s_{0}$, policy updates $L$, sample size $M$, policy and Q-network architectures, learning rate $\alpha$ for policy update, exploration scale $\sigma_{\rm{explore}}$;
@@
7:   Sample design $\xi_{k}^{(m)}=\mu_{w}(k,s_{k}^{(m)})+\epsilon_{\text{explore}}$ and $y_{k}^{(m)}\sim p(\cdot\mid\theta^{(m)},\xi_{k}^{(m)},I_{k}^{(m)})$;
8:   Update state $s_{k+1}^{(m)}$, and compare $r_{T}^{S}(s^{(m)}_{k+1})$ with the learned continuation value $Q_{\eta}^{\pi,\psi}(k+1,s^{(m)}_{k+1},\mu_{w}(k+1,s^{(m)}_{k+1}))$ for stopping decision;
@@ 3.3 Training instabilities and curriculum learning
Although the PG method provides a principled framework, direct training can be unstable because stopping and continuation values are mutually dependent. The optimal stopping set $\mathcal{T}_{k}$ in (14) depends on the continuation value $Q_{k}^{\pi,\psi}(s_{k},\mu_{k}(s_{k}))$, which is itself defined under the same design and stopping policies $(\pi,\psi)$. Consequently, the stopping policy requires a Q-function computed under that very policy, inducing a fixed-point relationship $\psi^{*}=f(\pi,\psi^{*})$.
@@
In practice, this circularity manifests during joint training of the policy network (parameters $w$) and the Q-network (parameters $\eta$). Early in training, the design policy is inevitably inaccurate, producing weak experiments with limited information gain. The Q-network then learns pessimistic continuation values, which induce premature stopping and truncate trajectories before informative designs can be discovered. When the reward landscape contains multiple attraction basins associated with different stopping stages, this feedback can bias learning toward an early-stopping basin before the policy has explored trajectories that lead to higher long-horizon rewards. As a result, the circular dependency can trap training in a suboptimal early-stage policy and cause it to miss a better, or even globally optimal, design–stopping strategy. This issue is particularly pronounced in problems with strong sequential dependencies, where early stopping disrupts the learning of long-horizon experimental strategies.
@@
To overcome this challenge, we adopt a simple form of curriculum learning, originally introduced by Bengio et al. (2009), in which an algorithm is gradually exposed to increasingly difficult aspects of the task. We implement this through a stopping probability schedule $p_{\text{stop}}(\ell)$, where $\ell$ indexes training iterations. When the stopping condition $s_{k}\in\mathcal{T}_{k}$ is satisfied, the algorithm follows the optimal stopping rule with probability $p_{\text{stop}}(\ell)$, and overrides it (forcing continuation) with probability $(1-p_{\text{stop}}(\ell))$ during training only. This strategy deliberately relaxes stopping early in training, producing longer trajectories that allow both the policy and Q-networks to improve before stopping decisions dominate the data distribution. As training progresses, $p_{\text{stop}}(\ell)$ is gradually increased, ensuring that the learned policies converge to the optimal stopping fixed point. In this way, curriculum learning breaks the self-reinforcing cycle of premature stopping and stabilizes training, particularly in settings with strong sequential dependencies.
In Section 4, we empirically examine when this curriculum strategy is beneficial. The clearest gains arise in a movement-constrained source detection problem, where premature stopping prevents the policy from learning informative later-stage designs.
3.4 Computational Considerations
The policy gradient method relies on Monte Carlo trajectory simulation during training to estimate gradients. The primary computational costs arise from evaluating information-theoretic rewards (KL divergences between posterior distributions) and forward model evaluations. For low-dimensional parameter spaces ($N_{\theta}\leq 4$), we employ grid-based discretization to compute posteriors and KL divergences, which is both efficient and transparent. For higher-dimensional problems, practitioners can instead leverage well-established approximation techniques, such as variational inference (Foster et al., 2019), MCMC-based density estimation, or amortized posterior inference via diffusion models (Baldassari et al., 2023), depending on problem structure and available computational resources.
Once trained, the learned policies enable computationally efficient decision-making at deployment. Determining the next design and stopping decision requires only a single forward pass through the policy and Q-networks, together with one posterior update and forward model evaluation. This eliminates the need for expensive online planning or repeated forward simulations, making the approach well suited for real-time sequential experimentation, even in resource-constrained settings.
@@ 4 Numerical Experiments
We evaluate the proposed framework on three numerical examples designed to test increasingly difficult aspects of the method. The linear-Gaussian benchmark validates the policy gradient algorithm against analytical solutions. The one-dimensional nonlinear problem tests non-conjugate Bayesian updating while retaining transparent posterior computation. The contaminant source detection problem introduces constrained sensor movement and stronger sequential dependence, making it a stress test for the curriculum strategy. Training hyperparameters are reported in Appendix A.5; they largely follow Shen and Huan (2023). Further hyperparameter tuning may improve performance but is not pursued here.
In our implementation, we adopt an adaptive sigmoid-based schedule to control the gradual increase of $p_{\text{stop}}(\ell)$. Specifically, we apply a shifted and scaled sigmoid function of the form:
$p_{\text{stop}}(\ell)=\frac{1}{1+\exp(-a(\ell-\ell_{0}))}$ (25)
where $a$ controls the transition steepness and $\ell_{0}$ determines the midpoint. We set parameters such that $p_{\text{stop}}(\ell)$ starts near zero, increases smoothly through mid-training, and saturates above 0.999 for the final 30 iterations to guarantee convergence to the optimal stopping policy. This sigmoid-based schedule provides a smooth transition that balances exploration during early training with stability near convergence, avoiding the abrupt changes that can destabilize actor-critic learning.
@@ 4.1 Linear-Gaussian Benchmark
with prior $\theta\sim\mathcal{N}(0,3^{2})$. Experiments are constrained to designs $\xi_{k}\in[0.1,3]$. Conjugacy enables closed-form characterization of optimal stopping and design policies (see Appendix A.4 for detailed derivation). We list the optimal utilities for varying number of experiments in Table 1. The optimal stopping stages are highlighted in red.
Table 1: Analytical optimal utility under varying experimental horizons and costs for the linear-Gaussian benchmark.
NN 1 2 3 4
$U(\Xi_{N})(c_{k}=0)$ 2.203 2.547 2.749 2.892
$U(\Xi_{N})(c_{k}=-0.5)$ 1.703 1.547 1.249 0.892
$U(\Xi_{N})(c_{k}=-0.25)$ 1.953 2.047 1.999 1.892
@@
For zero cost ($c_{k}=0$), the analytical solution shows that stopping is always optimal at the horizon $N$, since the KL divergence reward is non-decreasing. Figure˜4 (left column) shows that our method converges to the horizon $N=3$, achieving the analytical optimum, whereas naïve threshold rules stop prematurely and underperform.
For negative costs, Figure 4 (middle and right columns) illustrates two representative regimes. In the $N=3$ problem with $c_{k}=-0.5$, both vanilla and curriculum methods converge to the optimal early stopping at stage 1, consistent with the analytical solution. In the $N=4$ problem with $c_{k}=-0.25$, both methods achieve near-optimal final performance but exhibit different convergence behavior. The vanilla approach often underestimates the optimal stopping stage, becoming trapped below the analytical solution, whereas the curriculum method converges more reliably by exploring longer trajectories before settling at the correct stage. These results suggest that curriculum learning can improve the robustness of stopping-stage convergence when vanilla training tends to stop too early, although the effect is modest in this benchmark. The small discrepancy from the analytical solution in the $N=4$ case reflects a flat reward landscape, where stopping at stages 1–3 yields nearly indistinguishable rewards. Even in this setting, the PG method captures the relevant trade-offs, validating the proposed optimization approach. The remaining gap is attributed primarily to neural-network approximation and optimization choices, such as architecture and learning rate.
The distributional analysis in Figure˜5 further confirms that the trained policy learns the analytical structure of the linear-Gaussian benchmark. The stopping-stage histogram shows that the policy predominantly terminates at stage 3 when experimental costs are absent. The design histogram shows that designs concentrate near the upper bound ($\xi=3$), where the signal-to-noise ratio is maximized. Together, these distributions indicate that the learned policy recovers both the optimal stopping behavior and the optimal design behavior in this benchmark.
4.2 One-dimensional nonlinear test problem
To further assess the method beyond the conjugate setting, we consider a one-dimensional nonlinear test problem. The forward model is
$y_{k}=G(\theta,\xi_{k})+\epsilon_{k}=\theta^{3}\xi_{k}^{2}+\theta\exp(-|0.2-\xi_{k}|)+\epsilon_{k},\qquad\epsilon_{k}\sim\mathcal{N}(0,0.03^{2}),$ (27)
with prior $\theta\sim\mathcal{U}(0,1)$ and design constraint $\xi_{k}\in[0,1]$. Unlike the preceding linear-Gaussian benchmark, this model is non-conjugate: the posterior is no longer Gaussian, and the expected continuation value depends on the realized observations through the full posterior shape. At the same time, the scalar parameter keeps posterior inference accurate and transparent, allowing us to isolate the effect of nonlinearity without adding unnecessary computational complexity.